‏Ponszko Nursery used high - low data from June and July to determine its variable cost of 518 per unit . Additional information follows : Month Units produced Total.costs June 2.000 $ 48.000 July 1,000 30,000 If Ponszko's produces 2,300 units in August , how much is its total cost expected to be ? Format $ 552,000 Answer :


1
Expert's answer
2021-05-03T10:51:33-0400

"48000 - (518\times 2,000) =-988 000"


"(2 300 \times 518) + (-988 000) = 203 400"


$ 203, 400